OpenStudy (anonymous):

Describe the nature of the roots for this equation. x^2+3x-7=0 A. Two real, rational roots B. Two real, irrational roots C. Two complex roots D. One real, double root

OpenStudy (helder_edwin):

in an equation like \[ ax^2+bx+c=0 \] the solutions are given by \[ \LARGE x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \] depending on the value of \[ \Delta=b^2-4ac \] you have either (i) two real roots (ii) one (repeated) real root (iii) two complex roots

OpenStudy (campbell_st):

use b^2 - 4ac 3^2 - 4*1*-7 = 33 then there are 2 unequal irrational roots